# Broken Expectations: The Effects of Modelling Assumptions on the Inferred Dark Matter Distribution in the Milky Way’s Satellites

### v1

@@ 2.1.2 Initial Conditions
Following the approach of Iorio et al. (2019), we use an iteration procedure to converge on the initial properties of the dark matter halo and the stars, such that the stellar population has similar photometric and kinematic properties as observed in the dSphs today. We discuss this procedure further in Section 2.1.4 and, in the following, we describe the initial guesses for the dark matter halo and stellar distribution parameters and how the dwarfs are set on an orbit around the Milky Way potential.
@@
For the initial masses of the dark matter haloes, we take the estimates from Read & Erkal (2019), who determine pre-infall virial masses through abundance matching with the mean star formation rate. We then use the mass-concentration-redshift relation in CDM of Ludlow et al. (2016) to obtain the concentration of the NFW halos222We take the concentration at z=0 rather than at the redshift of infall as the relation of Read & Erkal (2019) is calibrated on inferred halo masses of isolated dwarfs at the present day. In any case, the concentration values will change as we use the iteration procedure to determine dark matter halo parameters that best fit the data in Section 2.1.4. . The stellar masses for the 5 dwarfs we investigate are taken from de Boer et al. (2014), de Boer et al. (2012a), de Boer et al. (2012b), Carrera et al. (2002), and Aparicio et al. (2001). As for the Plummer profile, we use the data stored within the GravSphere framework333 https://github.com/justinread/gravsphere for stellar kinematic and photometric data, which they obtain from observations made by Walker et al. (2009a), Spencer et al. (2018), and Walker et al. (2015). Using this data, we fit normalised Plummer mass profiles to obtain the Plummer scale radius , which we input in MAGI.
We use the virial and tidal radii as the exponential truncation radius for the NFW halo and as the cutoff for the Plummer profile, respectively444This choice is motivated by the fact that the outermost particles are quickly removed once the dwarf is placed in a Milky Way potential, and therefore this choice allows us to keep the resolution within the dwarf galaxy.. For the tidal radius, we use the derivation in Binney & Tremaine (2008):
@@
To set the dwarf spheroidals in orbit, we use galpy (Bovy, 2015) to obtain the 3-D positions and velocities of each dwarf in the ‘light’ and ‘heavy’ Milky Way potentials as a function of time. We then determine the infall time of each dwarf as the orbital apocenter nearest to the infall time estimates of Rocha et al. (2012)555This was determined through the deduction of a correlation between orbital energy and the time at which dSphs last cross the Milky Way virial radius, found through cosmological simulations.. Although this infall time estimate is derived from cosmological simulations, it nevertheless does not select Milky Way analogues with an LMC companion and thus our simulations are unlikely to represent the full tidal history of the dwarf sample. The position and velocity vectors at this point are used to initialise the orbit of each dwarf. We list the relevant orbital parameters in Table 1.
It should be noted that after initialisation with MAGI, the dSphs were left to simulate with GADGET-4 for 5 Gyr without an external potential for them to settle in equilibrium and minimize any numerical artefacts. Only thereafter were the dSphs ‘shifted’ to the starting apocenter radius, had their velocities ‘boosted’ into orbit, and left to evolve in heavy and light Milky Way potentials until their positions at the present day. We additionally note that since velocity anisotropy (isotropic) is disfavoured for cored stellar profiles living in cuspy dark matter haloes (An & Evans, 2009; Sánchez Almeida et al., 2023), the resultant stellar systems instead typically show radially varying velocity anisotropy, with a preference for mildly tangential anisotropy in the centre.
@@ 2.1.3 Evolution
We utilize the -body code GADGET-4 (Springel et al., 2021) to simulate the evolution of the dSphs. In GADGET-4, there are two methods for calculating gravitational interactions; however, for this study, we use the one-sided classical tree algorithm. The particle-particle interactions are softened with a spline kernel with a Plummer-equivalent softening length , with interactions outside 2.8 being exactly Newtonian. For time integration, we use the standard nested leapfrog time integration scheme with a hierarchy of time bins. Time steps are computed based on the absolute acceleration of a particle , the softening length , and a specified dimensionless accuracy parameter , such that:
where we use the default value . The time steps of each particle are then binned in time bins based on time durations that divide the current simulation time by powers of two.
Apart from these functionalities, GADGET-4 also offers group finding methods, which use the classic friends-of-friends (FOF) approach to find groups of particles of approximately virial overdensity (Springel et al., 2001a), and the SUBFIND algorithm to identify gravitationally bound substructures in these groups in configuration space (Springel et al., 2001b; Dolag et al., 2009). We use this to identify which particles get tidally stripped during the simulation.
Although GADGET-4 also includes an external potential (in our case, that of the Milky Way), it only offers a Hernquist potential. Thus, we modify GADGET-4 to read from a table of accelerations binned in cylindrical coordinates across the simulation volume. The acceleration at a specific coordinate is then calculated from linear interpolation between bins. This method was adapted from Tress et al. (2020) used in AREPO (Springel, 2010). We use galpy to construct the table of accelerations for the ‘light’ and ‘heavy’ Milky Way potential models.
@@ 2.1.4 Dwarf Spheroidal Galaxy Sample
In order to produce snapshots of dSphs that match their stellar masses and dynamics from observations, we use an iterative procedure from Iorio et al. (2019) to adjust the initial virial masses, stellar masses, NFW scale radii, and Plummer scale radii. Table 3 summarises the information used to initialise each dSph with MAGI. Figure 1 shows the resulting virial mass and concentration compared to the concentration-mass relation in CDM of Ludlow et al. (2016). Interestingly, many of the simulated dwarfs, particularly in the ‘heavy’ Milky Way potential, are outliers in the mass-concentration relation (although we note that the 0.1 dex scatter (Dutton & Macciò, 2014) we show here may not be representative of scatter at small halo masses, as seen in Ludlow et al. 2016). This may suggest that the infall times of Rocha et al. (2012) occur too early for these dwarfs or that the mass of the Milky Way must be significantly lower than for dwarf galaxy properties to be consistent with CDM. This could be an interesting avenue to explore for studies aiming to constrain the mass of the Milky Way. As we will show in Section 3.1.3, the line-of-sight velocity dispersion profiles of dwarf galaxies may also be indicative of the Milky Way’s mass.
In Figure 2, we show the resulting line-of-sight velocity dispersions within the projected half-light radius for each dwarf from simulated and real data from Walker et al. (2009a, 2015) and Spencer et al. (2018). Additionally, Figure 3 shows an example of the evolution of the line-of-sight velocity dispersion of Sculptor. It is clear that the iterative procedure worked well for all the dSphs except Ursa Minor, which could not converge to the real velocity dispersion despite numerous iterations in the light and heavy Milky Way potentials. This dSph is the most tidally affected in our sample, and due to the considerable tidal stripping, the Plummer parameters are adjusted simultaneously with the NFW parameters. Because the severity of the tidal stripping on the stellar component increases as the halo mass is reduced, this leads to an increase in suggested stellar mass, which in turn affects the line-of-sight velocity dispersion. This creates an unending feedback loop of mismatched parameters, causing the simulated line-of-sight velocity dispersion to never align with the observations. This does not affect the other dSphs since they lose a smaller fraction of stellar matter. This, therefore, suggests that the method from Iorio et al. (2019) may not be effective under conditions where the tidal stripping of the stellar component is severe.
@@
We have chosen quite a varied sample of dSphs in terms of mass and closeness to the Galactic centre and orbital eccentricity, as evident in Table 1. Carina, for example, has a nearly circular orbit relative to other dSphs. In contrast, Draco and Ursa Minor are strongly eccentric and approach close to the Galactic centre in the heavy Milky Way potential. Fornax is the most massive dSph in dark and stellar matter, while Carina is two orders of magnitudes smaller.
Lastly, we note that throughout this work, we place the “observer” at infinity from the direction of the Milky Way’s centre rather than at 8 kpc in the Galactic disk. We have made this choice as we explore various orbital phases of the dSphs in this work and would thus prefer to remove the ambiguity of Solar motion around the Galaxy and the uncertainties associated with the spatial extent of dwarf galaxies on the sky. Since we explore initially spherical NFW halos, which at the present day lie at least 76 kpc from Milky Way’s centre (corresponding to at most 6 degree offset from the Solar position), this choice does not significantly affect the constrained line-of-sight properties of our simulated dwarf galaxies or the results of this work.
@@ 2.2 Dark Matter Distribution Inference
For classical dSphs orbiting the Milky Way, large sets of kinematic and photometric data have already been obtained, though mostly limited to line-of-sight velocities and surface densities. Therefore, certain steps must be taken before estimating the masses from such data using the spherical Jeans equation. One is to recover the 3-dimensional density from the measured projected surface density . This can be done through an Abel transformation resulting in:
@@
With Equations 4, 5, 6 one thus has a system of equations to retrieve a total mass profile from a kinematic tracer population, through which one can assume a parameterization of a mass profile and anisotropy profile , and perform a best-fit for their parameters. In effect, the predicted and (both of which are dependent on the projected radius ) are compared with the kinematic and photometric data during the fit. However, since typically only line-of-sight information of the tracers is known, becomes poorly constrained. Thus, the Jeans dynamical modelling in its projected form suffers from the mass-anisotropy degeneracy.
Several methods have been explored to break this degeneracy. However, we use a method that exploits the fourth-order projected virial theorem (Merrifield & Kent, 1990; Richardson & Fairbairn, 2013a, b, 2014; Read & Steger, 2017) (see also Wardana et al. (2025) for fourth-order Jeans equations), which utilizes two new equations for the so-called virial shape parameters (VSP):
@@
This work performs such a fit for the mass profile of dSphs with line-of-sight data using a Python framework called pyGravSphere666https://github.com/AnnaGenina/pyGravSphere (Genina et al., 2020), which was adapted from the original GravSphere code (Read & Steger, 2017). It is based on the ensemble sampler EMCEE (Foreman-Mackey et al., 2013).
We use the broken power-law option for the density profile, as done in Read & Steger (2017). This uses five spatial bins defined as logarithmically spaced fractions of the projected half-light radius , with the values . Within each bin, the profile follows a power law defined by slopes with parameters . The complete profile is thus defined as:
@@ 2.3 Wolf Mass Estimator
Jeans analysis requires sufficiently large samples of spectroscopic and photometric data. For many dwarf satellites of the Milky Way, particularly the ultra-faints, such samples are often unavailable. For this reason, a simpler method is sometimes preferred, offering accurate estimates of dark matter mass enclosed within a given radius (Walker et al., 2009b) that could then justify the collection of more data and more accurate Jeans analysis. One such method is the Wolf mass estimator (Wolf et al., 2010) – a means for accurately estimating the enclosed mass at the half-light radius that is largely insensitive to the anisotropy parameter . The estimator requires only the line-of-sight velocity dispersion and the (projected) half-mass radius, which is often available even with small stellar data samples. In this work, we therefore also want to test the accuracy of this estimator for a sample of satellites at different stages of their orbit, which could then inform the application of the Wolf et al. (2010) mass estimator to other dwarfs.
@@
In the absence of full 3-dimensional information, one can still make further approximations to Equation 13, such as to assume that the luminosity weighted average of the line-of-sight velocity dispersion is approximately equal to the velocity dispersion at , , which also approximates the squared total velocity dispersion . It is also common to approximate this as the mean velocity dispersion inside the half-light radius (Campbell et al., 2017; Genina et al., 2018). Furthermore, one can approximate the radius to the half mass radius . For a Plummer profile, . Similarly, one can find a way to approximate the half-mass radius with the projected 2-dimensional half-mass radius , with . Thus, the ‘2-D Wolf’ estimator for the enclosed mass at the half-mass radius is:
To get the luminosity-weighted average line-of-sight dispersion, we use (Kowalczyk & Łokas, 2014):
@@
From these fitted profiles, we obtain the projected half-mass radius777For the Plummer profile, we use the best-fit scale radius a, as the tidal modification is usually not utilized in the fit and is not converged, sometimes resulting in diverging mass profiles. and the profiles are further used to compute the luminosity-weighted line-of-sight velocity dispersion via Equation 16.
Both the ‘Wolf analysis’ and dynamical modelling with pyGravSphere are performed exclusively on the bound particles, as identified with SUBFIND. Any particles that are tidally stripped (identified as unbound) in any given snapshot are ignored in the analysis of future snapshots. In modelling with pyGravSphere, we pay additional attention to the line-of-sight velocity dispersion profiles, ensuring that no data points lie outside the tidal radius of the galaxy and are not significant outliers from the main body of the galaxy.

### v2

@@ 2.1.2 Initial Conditions
Following the approach of Battaglia et al. (2015) and Iorio et al. (2019), who produced models of Fornax and Sculptor orbiting in the Milky Way potential in a setup similar to ours, we use an iteration procedure to converge on the initial properties of the dark matter halo and the stars, such that the stellar population has similar photometric and kinematic properties as observed in the dSphs today. We discuss this procedure further in Section 2.1.4 and, in the following, we describe the initial guesses for the dark matter halo and stellar distribution parameters and how the dwarfs are set on an orbit around the Milky Way potential.
@@
For the initial masses of the dark matter haloes, we take the estimates from Read & Erkal (2019), who determine pre-infall virial masses through abundance matching with the mean star formation rate. We then use the mass-concentration-redshift relation in CDM of Ludlow et al. (2016) to obtain the concentration of the NFW halos444We take the concentration at z=0 rather than at the redshift of infall as the relation of Read & Erkal (2019) is calibrated on inferred halo masses of isolated dwarfs at the present day. In any case, the concentration values will change as we use the iteration procedure to determine dark matter halo parameters that best fit the data in Section 2.1.4. . The stellar masses for the 5 dwarfs we investigate are taken from de Boer et al. (2014), de Boer et al. (2012a), de Boer et al. (2012b), Carrera et al. (2002), and Aparicio et al. (2001). As for the Plummer profile, we use the photometric data stored within the GravSphere framework555https://github.com/justinread/gravsphere from McMonigal et al. (2014); Koposov et al. (2014) and Flewelling et al. 2020 (the data selection is described in section 4.2 of Read et al. 2019). Using this data, we fit normalised Plummer mass profiles to obtain the Plummer scale radius , which we input in MAGI.
We use the virial and tidal radii as the exponential truncation radius for the NFW halo and as the cutoff for the Plummer profile, respectively666This choice is motivated by the fact that the outermost particles are quickly removed once the dwarf is placed in a Milky Way potential, and therefore this choice allows us to keep the resolution within the dwarf galaxy.. For the tidal radius, we use the derivation in Binney & Tremaine (2008):
@@
To set the dwarf spheroidals in orbit, we use galpy (Bovy, 2015) to obtain the 3-D positions and velocities of each dwarf in the ‘light’ and ‘heavy’ Milky Way potentials as a function of time. We then determine the infall time of each dwarf as the orbital apocenter nearest to the infall time estimates of Rocha et al. (2012)777This was determined through the deduction of a correlation between orbital energy and the time at which dSphs last cross the Milky Way virial radius, found through cosmological simulations.. Although this infall time estimate is derived from cosmological simulations, it nevertheless does not select Milky Way analogues with an LMC companion and thus our simulations are unlikely to represent the full tidal history of the dwarf sample. The position and velocity vectors at this point are used to initialise the orbit of each dwarf. We list the relevant orbital parameters in Table 1.
It should be noted that after initialisation with MAGI, the dSphs were left to simulate with GADGET-4 for 5 Gyr without an external potential for them to settle in equilibrium and minimize any numerical artefacts. Only thereafter were the dSphs ‘shifted’ to the starting apocenter radius, had their velocities ‘boosted’ into orbit, and left to evolve in heavy and light Milky Way potentials until their positions at the present day. We additionally note that since velocity anisotropy (isotropic) is disfavoured for cored stellar profiles living in cuspy dark matter haloes (An & Evans, 2009; Sánchez Almeida et al., 2023), the resultant stellar systems instead typically show radially varying velocity anisotropy, with a preference for mildly tangential anisotropy in the centre. We note that Wolf & Bullock (2012), using Jeans dynamical modeling with constant anisotropy, also find that Milky Way’s dwarfs require slightly tangential anisotropies given their dark matter haloes have an NFW form.
@@ 2.1.3 Evolution
We utilize the -body code GADGET-4 (Springel et al., 2021) to simulate the evolution of the dSphs. GADGET-4 also offers group finding methods, which use the classic friends-of-friends (FOF) approach to find groups of particles of approximately virial overdensity (Springel et al., 2001a), and the SUBFIND algorithm to identify gravitationally bound substructures in these groups in configuration space (Springel et al., 2001b; Dolag et al., 2009). We use this to identify which particles get tidally stripped during the simulation and which remain bound.
To model the Milky Way potential, we modify GADGET-4 to read from a table of accelerations binned in cylindrical coordinates across the simulation volume. The acceleration at a specific coordinate is then calculated from linear interpolation between bins. This method was adapted from that of Tress et al. (2020) within the AREPO code (Springel, 2010). We use galpy to construct the table of accelerations for the ‘light’ and ‘heavy’ Milky Way potential models.
@@ 2.1.4 Dwarf Spheroidal Galaxy Sample
In order to produce snapshots of dSphs that match their stellar masses and dynamics from observations, we use an iterative procedure from Battaglia et al. (2015) and Iorio et al. (2019) to adjust the initial virial masses, stellar masses, NFW scale radii, and Plummer scale radii. Table 3 summarises the information used to initialise each dSph with MAGI. The top panel of Figure 1 shows the resulting virial mass and concentration compared to the concentration-mass relation in CDM of Ludlow et al. (2016). Interestingly, many of the simulated dwarfs, particularly in the ‘heavy’ Milky Way potential, are outliers in the mass-concentration relation (although we note that the 0.1 dex scatter (Dutton & Macciò, 2014) we show here may not be representative of scatter at small halo masses, as seen in Ludlow et al. 2016). This may suggest that the infall times of Rocha et al. (2012) occur too early for these dwarfs or that the mass of the Milky Way must be significantly lower than for dwarf galaxy properties to be consistent with CDM. This could be an interesting avenue to explore for studies aiming to constrain the mass of the Milky Way. As we will show in Section 3.1.3, the line-of-sight velocity dispersion profiles of dwarf galaxies may also be indicative of the Milky Way’s mass.
The dashed cyan line and shaded band show the mass-concentration relation from the simulations of Wang et al. (2020b), who explore 20 orders of magnitude in halo mass, covering the range of expected halo masses of our simulated dwarfs. It can be seen that the scatter in this relation is substantially larger, easing the tension with the dwarfs in the ‘heavy’ Milky Way potential, while ‘light’ Fornax remains an outlier. We note, however, that the simulations of Wang et al. (2020b) focus on the void regions of the CDM universe and thus may not be applicable to the expected concentrations of dwarf galaxies in Local Group-like environments (Hellwing et al., 2021).
The bottom panel of 1 shows the relation between the concentration parameter,
and from Moliné et al. (2023), with the maximum circular velocity occurring at the radius and the Hubble parameter at a given redshift . The relation changes as a function of , where is the distance of the subhalo from the host galaxy and is the virial radius of the host. The black error bar, shows the typical scatter in the relation in this range of for a given from Massari et al. (2020). Instead of individual values, we show the relation marginalised for ranges of xsub for clarity. In agreement with the mass-concentration relation in the upper panel, Draco and Ursa Minor in the ‘heavy’ potential, as well as Fornax in the ‘light’ potential are outliers in the relation, independent of .
@@
In Figure 3 and 4, we show the resulting line-of-sight velocity dispersion and surface density profiles for each dwarf from our simulations and real data from Walker et al. (2009a, 2015); Spencer et al. (2018); McMonigal et al. (2014); Koposov et al. (2014); Flewelling et al. (2020). We note that throughout this work, we place the “observer” at infinity from the direction of the Milky Way’s centre rather than at 8 kpc in the Galactic disk. We have made this choice as we explore various orbital phases of the dSphs in this work, with the dynamics and tidal stripping determined primarily by the motion of the dwarf with respect to the Galactic centre. We thus prefer to remove the ambiguity of Solar motion around the Galaxy and the uncertainties associated with the spatial extent of dwarf galaxies on the sky. Since we explore initially spherical NFW halos, which at the present day lie at least 76 kpc from Milky Way’s centre (corresponding to at most 6 degree offset from the Solar position), this choice does not significantly affect the constrained line-of-sight properties of our simulated dwarf galaxies or the results of this work, which we have also explicitly verified.
In the surface density, our simulated dwarfs are in good agreement with the observational data, particularly in the ‘light’ potential, with a notable exception of Carina, where the observational data show a nearly flat surface density in the central 200 pc, which is not well described by a Plummer profile (purple line). The Fornax ‘heavy’ model also overestimates the surface density within the central 300 pc. Fornax is known to not be well-fit by a single Plummer profile, since the outer profile falls off more steeply than allowed by the model (Moskowitz & Walker, 2020). In Ursa Minor, a single Plummer model also struggles to describe the profile shape, although our ‘light’ model is in good agreement with the observed surface density.
In the line-of-sight velocity dispersion, the agreement is again better in the ‘light’ than the ‘heavy’ potential and varies as a function of radius. For Fornax, the velocity dispersion in both cases is overestimated in the central 300 pc, while the value within the half-light radius is in good agreement between the model and observations.
From the agreement between the half-light radius and the velocity dispersion calculated within, it is clear that the iterative procedure worked well for all the dSphs except Ursa Minor, which could not converge to the observed velocity dispersion despite numerous iterations in the light and heavy Milky Way potentials. This dSph is the most tidally affected in our sample and, as pointed out by Battaglia et al. (2015), in the regime where the structure of the dwarf changes significantly due to tidal stripping and tidal shocks, the iterative algorithm may not be able to converge. This is because the resultant line-of-sight velocity dispersion and the half-light radius may not scale linearly with the changes in the dark matter halo and stellar parameters in each new parameter iteration. Convergence in this case is also likely affected by the stochasticity in each newly generated set of initial conditions, as the number of stellar particles is particularly low ().
Battaglia et al. (2015) modelled Fornax using a cored dark matter halo and an approximate deprojection of a Sersic (1968) surface density profile for the stars. Their model of surface density profile provided a better fit in the central regions of Fornax, which flattens towards the centre more quickly than the Plummer model; however, their recovered line-of-sight velocity dispersion profiles are comparable to ours, with difficulty in reproducing the lower velocity dispersion in the central 0.5 kpc.
Iorio et al. (2019) modelled Sculptor with a Plummer law, embedded in a cored NFW-like halo. Their models seem to slightly overestimate the surface density in the inner regions (< 100 pc) and underestimate it in the outer regions, while our models slightly overestimate the surface density around 50 pc. The line-of-sight velocity dispersion profiles in their models provide a similar quality of fit to the data as we find here.
@@ 2.2 Dark Matter Distribution Inference
For classical dSphs orbiting the Milky Way, large sets of kinematic and photometric data have already been obtained, though mostly limited to line-of-sight velocities and surface densities. It is possible to recover the 3-dimensional density from the measured projected surface density . This can be done through an Abel transform:
@@
With Equations 4, 5, 6 one thus has a system of equations to retrieve a total mass profile from a kinematic tracer population, through which one can assume a parameterization of a mass profile and anisotropy profile , and perform a best-fit for their parameters. In effect, the predicted and (both of which are dependent on the projected radius ) are compared with the kinematic and photometric data during the fit. However, since typically only line-of-sight information of the tracers is known, becomes poorly constrained. Thus, the Jeans dynamical modelling suffers from the mass-anisotropy degeneracy.
Several methods have been explored to break this degeneracy. However, we use a method that exploits the fourth-order projected virial theorem (Merrifield & Kent 1990; Richardson & Fairbairn 2013a, b, 2014; Read & Steger 2017; see also Wardana et al. (2025) for fourth-order Jeans equations), which utilizes two new equations for the so-called virial shape parameters (VSP):
@@
This work performs such a fit for the mass profile of dSphs with line-of-sight data using a Python framework called pyGravSphere888https://github.com/AnnaGenina/pyGravSphere (Genina et al., 2020), which was adapted from the original GravSphere code (Read & Steger, 2017). It is based on the ensemble sampler EMCEE (Foreman-Mackey et al., 2013).
We use the broken power-law option for the density profile, as done in Read & Steger (2017), Read et al. (2018) and Read et al. (2019). This uses five spatial bins defined as logarithmically spaced fractions of the projected half-light radius , with the values , where is fixed to the scale radius of a single Plummer profile fit. Within each bin, the profile follows a power law defined by slopes with parameters . The complete profile is thus defined as:
@@
Finally, we point out that unlike the works of Battaglia et al. (2015) and Iorio et al. (2019), we chose to carry out Jeans analysis on the kinematic sample of bound stars as determined by SUBFIND, as opposed to a more observationally motivated kinematic sample selection. This is because we aim to study the departure of the orbiting dwarfs from dynamical equilibrium as a result of tides, as opposed to the effects of contamination of kinematic samples by tidally stripped stars, which has been previously explored in the literature (Klimentowski et al., 2007, 2009; Łokas et al., 2010); however, as we make our dwarf models publicly available it would be straightforward to incorporate these effects in future investigations.
@@ 2.3 Wolf Mass Estimator
For many dwarf satellites of the Milky Way, particularly the ultra-faints, large spectroscopic samples are often unavailable. In these cases, accurate estimates of mass enclosed within the half-light radius can be obtained with simple estimators (Walker et al., 2009b). One such method is the Wolf mass estimator (Wolf et al., 2010), which is largely insensitive to the anisotropy parameter . The estimator requires only the line-of-sight velocity dispersion and the (projected) half-mass radius, which is often available even with small stellar data samples. In this work, we therefore also want to test the accuracy of this estimator for a sample of satellites at different stages of their orbit, which could then inform the application of the Wolf et al. (2010) mass estimator to other dwarfs.
@@
In the absence of full 3-dimensional information, one can still make further approximations to Equation 13 to obtain the ‘2-D Wolf’ estimator:
with the luminosity-weighted average line-of-sight dispersion (Kowalczyk & Łokas, 2014):
@@
From these fitted profiles, we obtain the projected half-mass radius999For the Plummer profile, we use the best-fit scale radius , as the tidal modification is usually not utilized in the fit and is not converged, sometimes resulting in diverging mass profiles. and the profiles are further used to compute the luminosity-weighted line-of-sight velocity dispersion via Equation 16.
Both the ‘Wolf analysis’ and dynamical modelling with pyGravSphere are performed exclusively on the bound particles, as identified with SUBFIND. Any particles that are tidally stripped (identified as unbound) in any given snapshot are ignored in the analysis of future snapshots. Aside from these constraints, we use all bound particles in our analysis. We note that this at least doubles the spectroscopic sample used in Read et al. (2019) for poorer resolved systems like Ursa Minor, but is comparable to samples expected in the near future from e.g. Prime Focus Spectrograph on Subaru (Kirby & PFS Galactic Archaeology Team, 2021). In modelling with pyGravSphere, where the velocity dispersion profiles can be particularly finely binned, with binning, we pay additional attention to the line-of-sight velocity dispersion profiles, ensuring that no data points lie outside the tidal radius of the galaxy and are not significant outliers from the main body of the galaxy.